OpenStudy (anonymous):

Write an equation of the line that passes through point P and is parallel to the line with the given equation: P(-2,0); 2y=-x+6

OpenStudy (anonymous):

Since it is parallel, the slope will be the same in both equations, so slope = m = -1 then use the point slope formula y-y1=m(x-x1) to figure out the equation with the (x1,y1) point you are given

Directrix (directrix):

2y=-x+6 y = (-1/2) x + 3 so the slope of the parallel line is also (-1/2) y - 0 = (-1/2)( x - (-2)) y = (-1/2)( x + 2) y = (-1/2) x - 1 Click on the link to look at the two graphs: http://www.wolframalpha.com/input/?i=%28y+%3D+%28-1%2F2%29+x+-+1%29+and+%282y%3D-x%2B6%29